OpenStudy (anonymous):

How would i find the simplified version of 3 log x – log y + log z ?

OpenStudy (anonymous):

3log(x*z/y)

OpenStudy (anonymous):

\[3\frac{ \log (xz/y )}{ 1 }=\frac{ \log ((xz) ^3/y)}{ 1 }\]

OpenStudy (anonymous):

\[\log(\frac{ x^3z }{ y })\]

OpenStudy (anonymous):

would it be \[\log(\frac{ x^3z }{ y }) or \log(\frac{ x^3y }{z }) ?\]

OpenStudy (anonymous):

the first one

OpenStudy (campbell_st):

just to avoid @casanovaxx answers you need law 1\[alog(x) = \log(x^a)\] next you will need law 2 \[alog(x) + \log(z) = \log(x^az)\] law 3 \[\log(x^az) - \log(z) = \log(\frac{x^az}{y})\] maybe this avoids confusion

OpenStudy (anonymous):

thanks both!!

OpenStudy (anonymous):

;)